OpenStudy (anonymous):

in triangle RST , m< R = 104 degrees,r=20 , and t= 10. find the m

OpenStudy (amistre64):

drawing the triangle always helps; show me what you thing the triangle should look like

OpenStudy (anonymous):

i dont exactly know how to do the triangle

OpenStudy (anonymous):

i dont exactly know how to do the triangle

OpenStudy (amistre64):

you just draw up a general triangle, and label the parts, like the other one i did for you. I need you to at least try to draw it ... i will correct it if there is an error.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

 the angles should have been capitalized

OpenStudy (amistre64):

thats perfect :) and yes, the angles should look different than the sides to help avoid confustion; now lets label what we know so that we can develop a plan for what we dont know

OpenStudy (anonymous):

 ok

OpenStudy (anonymous):

we are working to find the measure of angle s

OpenStudy (amistre64):

we can find T by the law of sines: and thereby find the S since we know T and R

OpenStudy (anonymous):

how do we do that

OpenStudy (amistre64):

\[\frac{sinT}{t}=\frac{sinR}{r}\] \[sinT=\frac{t~sinR}{r}\] \[T=sin^{-1}\left(\frac{t~sinR}{r}\right)\]

OpenStudy (anonymous):

T= sin -^1 (10 (.9703)/20)

OpenStudy (anonymous):

sorry i wwas having complications with the computer

OpenStudy (amistre64):

that should simplify to something like: 29.02 degrees

OpenStudy (anonymous):

after doing that i got T= sin ^-1(.48515)

OpenStudy (anonymous):

and then i got 29.0223 so round to the nearest degree would be 29

OpenStudy (amistre64):

lets avoid rounding at this stage, we want to find angle S, this is just angle T

OpenStudy (anonymous):

but t isnt what we were looking for so now we use the information and find the measure of angle s correct?

OpenStudy (anonymous):

oh ok

OpenStudy (amistre64):

R + T + S = 180 S = 180 - (R + T) = 180 - (133.02)

OpenStudy (anonymous):

46.98=S

OpenStudy (amistre64):

yep, and rounded we get?

OpenStudy (anonymous):

47

OpenStudy (amistre64):

great job !

OpenStudy (anonymous):

thank you i have difficulty with the law of sines